# SFVInt: Simple, Fast and Generic Variable-Length Integer Decoding using Bit Manipulation Instructions

### v1

@@ Abstract.
The ubiquity of variable-length integers in data storage and communication necessitates efficient decoding techniques. In this paper, we present SFVInt, a simple and fast approach to decode the prevalent Little Endian Base-128 (LEB128) varints. Our approach, distilled into a mere 500 lines of code, effectively utilizes the Bit Manipulation Instruction Set 2 (BMI2) in modern Intel and AMD processors, achieving significant performance improvement while maintaining simplicity and avoiding overengineering. SFVInt, with its generic design, effectively processes both 32-bit and 64-bit unsigned integers using a unified code template, marking a significant leap forward in varint decoding efficiency. We thoroughly evaluate SFVInt’s performance across various datasets and scenarios, demonstrating that it achieves up to a 2x increase in decoding speed when compared to varint decoding methods used in established frameworks like Facebook Folly and Google Protobuf.
${}^{\dagger}$ Work performed at ByteDance Infrastructure System Lab
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
In data processing and storage, handling variable-length integers efficiently is of paramount importance. These integers feature prominently across numerous systems such as search engines like Apache Lucene (Foundation, 2011), databases including IBM DB2 (Bhattacharjee et al., 2009) and Apache Kudu (Lipcon et al., 2015), and popular file formats such as ORC (Foundation, 2013a) and Parquet (Foundation, 2013b), acting as a key component for compact data representation. Their significance extends to serialization frameworks, notably Google Protobuf (Google, 2001), and the WebAssembly binary encoding (Consortium, 2022); they are also an essential part of the Go programming language’s default API (Griesemer, 2011), highlighting its widespread adoption across various critical technological domains (Liao et al., 2023; Bhattacherjee et al., 2021; Liao and Abadi, 2023; Liao, 2022). The Little Endian Base-128 (LEB128) encoding (Wikipedia, 2023), used for these variable-length integers (varints), is widely adopted for its optimal balance of space efficiency and implementation simplicity. However, a critical challenge arises in varint decoding, particularly due to the unpredictable lengths of encoded integers, which often lead to frequent branch mispredictions and thus become a bottleneck in performance.
Recognizing this challenge, we turn our attention to the Bit Manipulation Instruction Set 2 (BMI2) (bmi, 2013), an advancement in processor instruction sets available in modern Intel (Intel Corporation, 2016) and AMD CPUs (AMD, 2023). BMI2 offers a suite of instructions that are particularly adept at handling bitwise operations, a core component of varint decoding processes. Leveraging these capabilities, we propose SFVInt, a generic, simple, and fast approach to accelerate the decoding of LEB128 varints. Condensed into around 500 lines of code, our approach significantly outperforms traditional methods, offering up to a 2x increase in decoding speed over techniques used in well-established systems and libraries. Our approach is both simple and highly effective, capable of handling 32-bit and 64-bit unsigned integers with a unified C++ template. We provide a detailed analysis, highlighting how the use of BMI2 instructions contributes to this substantial performance enhancement.
To the best of our knowledge, this is the first time the BMI2 technique has been applied in the context of varint decoding, marking a significant advancement in this field. This research presents a critical enhancement in the efficiency of varint decoding, opening pathways for more refined data processing optimizations. The ramifications of this discovery are extensive, potentially influencing the design and performance optimization of databases, network systems, and big data applications, crucial in a data-centric era.
@@ 3.1. Basic Varint Operations
Algorithm: LEB128 Integer Encoding
@@
If val is 128 (0x80) or greater, the function stores the lowest 7 bits of val into the buf, setting the 8th bit of this byte (denoted by 0x80) to signal that more bytes follow. The buffer pointer buf is then advanced, and val is right-shifted by 7 bits to prepare the next 7 bits for encoding. This process repeats until val is less than 128. At this point, val is small enough to fit into the final byte, indicating the end of the variable-length integer.
@@
Algorithm: LEB128 Integer Decoding (Basic Version)
@@
The decoding algorithm reads each byte, combining its lower 7 bits into the result. The most significant bit indicates if subsequent bytes are part of the integer. The decoding concludes when a byte with the most significant bit unset is encountered, indicating the final byte of the varint. In line 1, the maximum shift values for LEB128 decoding – 28 for 32-bit integers and 63 for 64-bit integers – are determined by the integer size and the 7-bit grouping used in this encoding. For example, for a 32-bit unsigned integer, which has a maximum value of 4,294,967,295 ($2^{32}$ 1), it would require up to 5 bytes for representation, as 32 divided by 7 is approximately 4.57.
Varint Skipping: In columnar storage systems, efficient skip operations are vital for effective filtering and scanning of varint-encoded data streams. We employ a technique that processes 64-bit data blocks, enabling simultaneous handling of multiple varints.
Algorithm: LEB128 Integer Skipping
@@
In line 3, applying the bitwise NOT operation ($\sim$) inverts all bits of a 64-bit integer, turning continuation bits from ’1’s to ’0’s and vice versa. With the mask 0x8080808080808080, highlighting MSBs in each 8-bit segment, the AND operation (&) then isolates these MSBs. Utilizing SIMD instructions, specifically POPCNT, this step efficiently identifies varint terminations within the 64-bit block.
Varint Sizing: In varint-based systems, accurately estimating storage needs is crucial. We utilize precomputed lookup tables (LUTs) to quickly calculate the exact byte size needed for varint encoding, based on the position of the most significant set bit in integers. This approach optimizes computational efficiency with low overhead in caching, facilitating rapid estimations for dynamic storage allocation.
Algorithm: LEB128 Integer Sizing
@@
6.       size += kLUT32[__builtin_clz  (val | 1)]
@@
8.       size += kLUT64[__builtin_clzll(val | 1)]
@@
This algorithm computes the total storage requirement for varint-encoded integers by iteratively applying the LZCNT/BSR instruction to count leading zeros, adjusted for zero values with (val|1). The outcome indexes into predefined lookup tables (LUTs) for 32-bit and 64-bit integers (kLUT32, kLUT64) to retrieve byte sizes.
@@ 3.2. BMI2-Enhanced Bulk Varint Decoding
Mask Configuration: The first step involves employing a 64-bit mask with the _pext_u64 BMI2 instruction for efficient extraction of most significant bits (MSBs) from byte sequences. Configurations like 0x0000808080808080 for six bytes, 0x0080808080808080 for seven bytes, and 0x8080808080808080 for eight, cater to different ranges of extracted values (0-63, 0-127 and 0-255, respectively), influencing the number of processed bytes. Our implementation adopts the six-byte mask 0x0000808080808080, based on findings that longer masks risk L1 instruction cache spillage, compromising performance. This choice ensures an optimized balance between decoding efficiency and CPU resource utilization, in line with the operational constraints of modern CPU architectures.
@@
This method showcases adaptability to varint structures, identified by mask results, reducing the need for conditional byte-by-byte checks that can cause branch mispredictions and pipeline stalls.
Cross-Boundary Cases: In addressing cross-boundary cases where integers span two or even three consecutive blocks, we adopt a refined approach crucial for accurate and complete decoding. Our approach employs two key variables: shift_bits and partial_value. shift_bits tracks the necessary bit displacement for integers crossing block boundaries, while partial_value retains the already decoded segment from a previous block. This system ensures proper alignment and integration of integer segments across blocks, merging partial_value with new segments to reconstruct complete integers.
@@
Algorithm: The following presents an advanced varint decoding algorithm optimized for bulk processing with a 6-byte MASK, integrating cross-boundary case handling and fully leveraging BMI2 instructions for improved efficiency and speed. Line 7 employs _pext_u64 to extract MSBs from a 64-bit word using a defined mask, revealing the varint structure, including the count and distribution of integers. The algorithm then uses this output (mval) in a switch-case to customize the decoding for each integer’s unique encoding pattern.
Algorithm: Optimized Bulk Varint Decoding with BMI2
@@
18:      case 15:
19:         cu_val = _pext_u64(word, 0x0000007f7f7f7f7f);
@@
21:         *res++ = _pext_u64(word, 0x00007f0000000000);
22:         shift_bits = 0; pt_val = 0; n -= 2;
23:      case 32:
24:         cu_val = _pext_u64(word, 0x000000000000007f);
25:         *res++ = (cu_val << shift_bits) | pt_val;
26:         *res++ = _pext_u64(word, 0x0000000000007f00);
27:         *res++ = _pext_u64(word, 0x00000000007f0000);
28:         *res++ = _pext_u64(word, 0x000000007f000000);
29:         *res++ = _pext_u64(word, 0x0000007f00000000);
30:         pt_val = _pext_u64(word, 0x00007f0000000000);
31:         shift_bits = 7; n -= 5;
32:      case 45:
33:         cu_val = _pext_u64(word, 0x0000000000007f7f);
34:         *res++ = (cu_val << shift_bits) | pt_val;
35:         *res++ = _pext_u64(word, 0x0000007f7f7f0000);
36:         pt_val = _pext_u64(word, 0x00007f0000000000);
37:         shift_bits = 7; n -= 2;
38:      case 62:
39:         cu_val = _pext_u64(word, 0x000000000000007f);
40:         *res++ = (cu_val << shift_bits) | pt_val;
41:         pt_val = _pext_u64(word, 0x00007f7f7f7f7f00);
42:         shift_bits = 35; n -= 1;
43:      case 63:
44:         pt_val |= _pext_u64(word, 0x00007f7f7f7f7f7f)
45:                << shift_bits;
46:         shift_bits += 42;
47:      case ... // Other cases omitted for brevity
48: end while // fallback later
@@
In an example from lines 38-42, the algorithm manages a case where a significant part of an integer starts in a preceding 64-bit block and continues into the current one. The binary pattern 62 (represented as 111110) under a 6-byte MASK indicates an integer’s end within the current block. Using _pext_u64, the algorithm merges this block’s decoded segment with the preceding block’s pt_val, completing the integer. Additionally, the presence of five leading 1s (11111) signals an incomplete integer extending into the next block, which is extracted and stored in pt_val for subsequent decoding, with shift_bits adjusted to 35 for alignment, i.e., the five 7-byte sequences. This decrement in n by 1 ensures the algorithm’s seamless management of cross-boundary integers, maintaining decoding continuity and precision.
Case 63, depicted in lines 44-45, addresses a binary representation of 111111, signaling an integer spanning three data blocks, with the current block as its mid-segment. This specific case requires extracting six bytes (42 bits) via _pext_u64, appending this data to the partial value to maintain integer continuity across blocks.

### v3

@@ Abstract.
The ubiquity of variable-length integers in data storage and communication necessitates efficient decoding techniques. In this paper, we present SFVInt, a simple and fast approach to decode the prevalent Little Endian Base-128 (LEB128) varints. Our approach effectively utilizes the Bit Manipulation Instruction Set 2 (BMI2) in modern Intel and AMD processors, achieving significant performance improvement while maintaining simplicity and avoiding overengineering. SFVInt, with its generic design, effectively processes both 32-bit and 64-bit unsigned integers using a unified code template, marking a significant leap forward in varint decoding efficiency. We thoroughly evaluate SFVInt’s performance across various datasets and scenarios, demonstrating that it achieves up to a 2x increase in decoding speed when compared to varint decoding methods used in established frameworks like Facebook Folly and Google Protobuf.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06journalyear: 2024copyright: rightsretainedconference: 20th International Workshop on Data Management on New Hardware ; June 10, 2024; Santiago, AA, Chilebooktitle: 20th International Workshop on Data Management on New Hardware (DaMoN ’24), June 10, 2024, Santiago, AA, Chiledoi: 10.1145/3662010.3663439isbn: 979-8-4007-0667-7/24/06
@@ 1. Introduction
Variable-length integers (varints) play a crucial role in optimizing space efficiency for integer data representation across a wide range of systems and applications. From search engines like Apache Lucene (Foundation, 2011) and databases such as IBM DB2 (Bhattacharjee et al., 2009) and Apache Kudu (Lipcon et al., 2015), to columnar formats like ORC (Foundation, 2013a) and Parquet (Foundation, 2013b), varints serve as a fundamental building block for compact data storage and processing. The importance of varints extends beyond traditional data management systems. They are extensively used in serialization frameworks such as Google Protobuf (Google, 2001) and the WebAssembly binary encoding (Consortium, 2022), highlighting their significance in cross-platform data exchange and web-based applications. Moreover, varints are an integral part of the Go programming language’s default API (Griesemer, 2011), further emphasizing their widespread adoption across various critical technological domains (Liao et al., 2023; Bhattacherjee et al., 2021; Liao and Abadi, 2023; Liao, 2022).
Among the various varint encoding schemes, the Little Endian Base 128 (LEB128) (Wikipedia, 2023) has emerged as a widely adopted standard due to its optimal balance of space efficiency and implementation simplicity. LEB128 encodes integers as a sequence of bytes, where the most significant bit of each byte indicates whether more bytes follow, allowing for variable-length representation. However, despite its space-saving advantages, the decoding process for LEB128 varints often becomes a performance bottleneck in data-intensive applications. The unpredictable lengths of the encoded integers lead to branch mispredictions and limit the opportunities for vectorization, hindering the overall efficiency of the decoding process.
To address this challenge, we leverage the Bit Manipulation Instruction Set 2 (BMI2) (bmi, 2013), a powerful set of instructions available in modern Intel (Intel Corporation, 2016) and AMD (AMD, 2023) CPUs. BMI2 offers advanced bit manipulation capabilities that can significantly accelerate the decoding of LEB128 varints. By harnessing these instructions, we introduce SFVInt, a simple, fast, and generic approach for varint decoding. SFVInt stands out for its simplicity and performance. With just 500 lines of code, our solution achieves up to a 2x speedup in decoding compared to widely-used systems and libraries. SFVInt’s generic design seamlessly handles both 32-bit and 64-bit unsigned integers using a unified C++ template, enhancing its versatility and maintainability. Our in-depth analysis highlights the critical role of BMI2 instructions in achieving this substantial performance improvement. We demonstrate how SFVInt strategically employs these instructions to extract and manipulate the relevant bits from the encoded data, enabling efficient decoding while minimizing branch mispredictions.
To the best of our knowledge, our work represents the first comprehensive evaluation of BMI2 instructions in the context of varint decoding, marking a significant advancement in this field. By leveraging the powerful bit manipulation capabilities offered by these instructions, we have achieved a critical enhancement in the efficiency of varint decoding. The implications of our research extend beyond the scope of this paper, potentially influencing the design and performance optimization of a wide range of data-intensive systems. Databases, network protocols, and big data frameworks are just a few examples of the domains that can benefit from the improved varint decoding performance brought forth by our approach.
@@ 3.1. Basic Varint Operations
Algorithm 1: LEB128 Integer Encoding
@@
If val is 128 (0x80) or greater, the function stores the lowest 7 bits of val into the buf, setting the 8th bit of this byte (denoted by 0x80) to signal that more bytes follow (line 2). The buffer pointer buf is then advanced, and val is right-shifted by 7 bits to prepare the next 7 bits for encoding (line 3). This process repeats until val is less than 128. At this point, val is small enough to fit into the final byte, indicating the end of the variable-length integer (line 5).
@@
Algorithm 2: LEB128 Integer Decoding (Basic Version)
@@
In the basic LEB128 decoding algorithm, each byte is read and its lower 7 bits are combined into the result integer using bitwise operations. The most significant bit of each byte serves as a continuation flag, indicating whether more bytes follow. The decoding process concludes when a byte with its most significant bit unset is encountered, signaling the end of the varint.
The maximum shift values for decoding are determined based on the target integer size (line 1). For 32-bit integers, the maximum shift is 28, while for 64-bit integers, it is 63. These values are derived from the fact that LEB128 uses a 7-bit encoding scheme. To illustrate this, consider a 32-bit unsigned integer with a maximum value of 4,294,967,295 ($12^{32}-$ 1). Representing this value using LEB128 would require up to 5 bytes, as $\lfloor 32/7\rfloor=4$, with 4 bits remaining. The shift values used in the decoding process correspond to each 7-bit group in these five bytes: 0, 7, 14, 21, and 28 (line 2).
This basic decoding algorithm serves as a foundation for understanding LEB128 decoding, but it may not be optimal for performance due to the byte-by-byte processing and the presence of conditional branches. More advanced techniques, such as vectorization and branch elimination, can be applied to improve decoding efficiency.
Varint Skipping: Efficient skip operations are crucial for effective filtering and scanning of varint-encoded data streams in columnar storage systems. By leveraging SIMD instructions and bitwise operations, we employ a technique that processes 64-bit data blocks, enabling simultaneous handling of multiple varints, improving performance compared to byte-by-byte approaches.
Algorithm 3: LEB128 Integer Skipping
@@
The algorithm above takes two input parameters: n, the number of integers to skip, and buf, a pointer to the buffer containing the encoded integers. It initializes a 64-bit pointer w to the same location as buf, enabling the processing of 64-bit words (line 1). The condition n >= 8 in line 2 ensures that the algorithm processes the buffer in 64-bit word-sized chunks as long as there are at least 8 integers to skip. This threshold is chosen because a 64-bit word can contain up to 8 bytes, each potentially representing a single-byte varint.
Within the loop, the algorithm performs a series of bitwise operations to determine the number of complete varints in the current 64-bit word (line 3). It first inverts all the bits of the dereferenced 64-bit word using the bitwise NOT operation ($\sim$). For LEB128 encoding, the most significant bit (MSB) of each byte being ’1’ indicates the continuation of a varint, and ’0’ indicates its termination. By inverting these bits, the continuation bits (’1’s) are turned into ’0’s and vice versa. Then, it applies a mask (0x8080808080808080) to isolate the MSBs of each byte, which indicate the continuation or termination of a varint. By performing the bitwise AND operation (&\&&) between the inverted word and the mask, the algorithm obtains a 64-bit value where the most significant bits of each byte are set to 1 if the corresponding varint terminates and 0 if it continues.
The POPCNT function is then used to efficiently count the number of set bits (varint terminations) in the resulting 64-bit value. This count represents the number of complete varints in the current word. The algorithm subtracts this count from n to update the number of remaining integers to skip (line 3). After the loop, the algorithm updates the buf pointer to the location where the 64-bit word processing stopped (line 5). It then enters a byte-by-byte skipping loop (lines 6-8) to handle any remaining integers. This loop continues until n reaches zero, indicating that the desired number of integers has been skipped.
By processing the buffer in 64-bit words and using efficient bitwise operations and SIMD instructions, the LEB128 Integer Skipping algorithm achieves fast skipping of varints, making it suitable for scenarios where large numbers of integers need to be skipped quickly.
Varint Sizing: In varint-based systems, accurately estimating storage needs is crucial for efficient memory allocation and management. We utilize a precomputed lookup table (LUT) to quickly calculate the exact byte size needed for varint encoding, based on the position of the most significant set bit in integers. This approach optimizes computational efficiency with low overhead in caching, facilitating rapid estimations for dynamic storage allocation.
Algorithm 4: LEB128 Integer Sizing
@@
6.       size += kLUT[__builtin_clz  (val | 1) + 32]
@@
8.       size += kLUT[__builtin_clzll(val | 1)]
@@
The algorithm computes the total byte size required for varint encoding a collection of integers. It leverages a precomputed LUT that maps the position of the most significant set bit in an integer to the corresponding varint byte size. The LUT is constructed based on the observation that the byte size needed for varint encoding depends on the magnitude of the integer value.
For each integer val in the input vector values, the algorithm determines the position of the most significant set bit using the LZCNT (Leading Zero Count) instruction, depending on the integer type (uint32_t or uint64_t). To handle the case where val is zero, the algorithm performs a bitwise OR operation with 1 (val|1) before counting the leading zeros [^1] It’s worth noting that BSF and BSR, with a source operand value of (0), leave the destination undefined and set the ZF (zero flag). However, C++20’s std::countl_zero doesn’t have this issue, if your compiler supports it.. This ensures that the LUT index is correctly calculated even for zero values. The position of the most significant set bit is then used as an index to retrieve the corresponding byte size from the LUT. For uint32_t values, an offset of 32 is added to the LUT index to account for the difference in bit positions between 32-bit and 64-bit integers.
To illustrate the algorithm 4, consider a small dataset of four 64-bit integers: 42, 1,337, 69,420, and 42,000,000. Figure 3 shows the binary representation of each integer, highlighting the most significant set bit position and the corresponding LUT index and byte size. The algorithm determines the varint encoding byte size for each integer using the precomputed LUT. For 42, the most significant set bit is at position 58, and the LUT value at index 58 is 1, indicating a 1-byte encoding. Similarly, 1,337 requires 2 bytes (LUT index 53), 69,420 requires 3 bytes (LUT index 47), and 42,000,000 requires 4 bytes (LUT index 38). The algorithm sums the byte sizes, yielding a total storage requirement of 10 bytes (1 + 2 + 3 + 4) for the dataset. The retrieved byte sizes for each integer are accumulated in the size variable, which represents the total storage requirement for the varint-encoded integers.
Figure 3. Varint Sizing Example
The use of a precomputed LUT eliminates the need for runtime calculations and conditional branching, resulting in efficient and fast estimation of varint sizes. The LUT is small enough to fit in the cache, minimizing cache misses and further improving performance. By employing this LUT-based approach, the algorithm achieves rapid and accurate estimation of storage requirements for varint encoding, enabling efficient memory allocation and management in varint-based systems.
@@ 3.2. BMI2-Enhanced Bulk Varint Decoding
Mask Configuration: The first step in our BMI2-based LEB128 varint decoding approach involves the use of a 64-bit mask with the _pext_u64 instruction for efficient extraction of the most significant bits (MSBs) from byte sequences. The choice of mask configuration is crucial to the performance of our approach, as it directly influences the number of bytes processed simultaneously and the range of values that can be extracted. In SFVInt, we considered three primary mask configurations: 0x0000808080808080 for six bytes, 0x0080808080808080 for seven bytes, and 0x8080808080808080 for eight bytes. Each of these configurations caters to different ranges of extracted values: 0-63 for six bytes, 0-127 for seven bytes, and 0-255 for eight bytes.
@@
This method showcases a high degree of adaptability to various varint structures, which are identified through the mask results obtained from the _pext_u64 instruction. By leveraging these mask results, our approach can efficiently process the encoded data stream while minimizing the need for conditional byte-by-byte checks, which are a common source of branch mispredictions and pipeline stalls in traditional decoding methods. This understanding of the mask’s role is fundamental to our optimized decoding process. By using the mask results to guide the decoding logic, we can efficiently extract and process the relevant bits from the encoded data stream, resulting in a significant performance improvement over traditional methods.
While longer masks, such as 0x8080808080808080, might appear to offer better performance by processing more bytes simultaneously, our empirical findings suggest otherwise. The use of longer masks results in a larger instruction footprint, as the decoding logic needs to handle a greater number of possible varint configurations. This increased complexity can exceed the capacity of the L1 instruction cache, leading to cache misses and performance degradation. To strike an optimal balance between decoding efficiency and CPU resource utilization, we empirically determined that the six-byte mask configuration provides the best performance. This choice aligns with the operational constraints of modern CPU architectures, minimizing the risk of L1 instruction cache spillage while still enabling efficient decoding.
Cross-Boundary Cases: In addressing cross-boundary cases where integers span two or even three consecutive blocks, we adopt a refined approach crucial for accurate and complete decoding. Our approach employs two key variables: shift_bits and partial_value. shift_bits tracks the necessary bit displacement for integers crossing block boundaries, while partial_value retains the already decoded segment from a previous block. This system ensures proper alignment and integration of integer segments across blocks, merging partial_value with new segments to reconstruct complete integers, thereby preserving the accuracy and integrity of the overall decoding process.
@@
Algorithm: The following presents an advanced varint decoding algorithm optimized for bulk processing with a 6-byte MASK, integrating cross-boundary case handling and fully leveraging BMI2 instructions for improved efficiency and speed. Line 7 employs _pext_u64 to extract MSBs from a 64-bit word using a defined mask, revealing the varint structure, including the count and distribution of integers. The algorithm then uses this output (mval) in a switch-case to tailor the decoding process for each integer, adapting to their specific encoding patterns within the data stream. To illustrate the decoding process, let’s examine three representative cases: case 0, case 45, case 62 and case 63.
Algorithm 5: Optimized Bulk Varint Decoding with BMI2
@@
18:      case 45:
19:         cu_val = _pext_u64(word, 0x0000000000007f7f);
@@
21:         *res++ = _pext_u64(word, 0x0000007f7f7f0000);
22:         pt_val = _pext_u64(word, 0x00007f0000000000);
23:         shift_bits = 7; n -= 2;
24:      case 62:
25:         cu_val = _pext_u64(word, 0x000000000000007f);
26:         *res++ = (cu_val << shift_bits) | pt_val;
27:         pt_val = _pext_u64(word, 0x00007f7f7f7f7f00);
28:         shift_bits = 35; n -= 1;
29:      case 63:
30:         pt_val |= _pext_u64(word, 0x00007f7f7f7f7f7f)
31:                << shift_bits;
32:         shift_bits += 42;
33:      case ... // Other cases omitted for brevity
34: end while
@@
Case 0 (lines 10-17) handles the scenario where the current 64-bit word contains six complete single-byte integers. When mval is 0, it indicates that all six bytes in the word represent individual integers. The algorithm proceeds to extract each integer using _pext_u64 with carefully crafted masks targeting their respective positions within the word. The extracted integers are then stored in the output array res, and the shift_bits and pt_val variables are reset to prepare for the next iteration. Finally, the count of remaining integers to decode (n) is decremented by 6 (line 17).
Case 45 (lines 19-23) handles the scenario where the current 64-bit word contains two 2-byte integers. When mval is 45 (binary representation: 101101), it indicates that the first integer spans across the first two bytes of the word, while the second integer spans across the third and fifth bytes. The algorithm extracts these integers using _pext_u64 with masks 0x0000000000007f7f and 0x0000007f7f7f0000, respectively (lines 19-21). The pt_val variable is updated with any remaining partial value (line 22), and the shift_bits variable is set to 7 to align the decoding process for the next iteration. Finally, the count of remaining integers (n) is decremented by 2 (line 23).
Case 62 (lines 25-28) demonstrates the algorithm’s ability to handle integers that span across multiple 64-bit words. When mval is 62 (binary representation: 111110) under a 6-byte mask, it signifies that an integer ends within the current word, with its remaining bits located in the preceding word. The algorithm extracts the relevant bits from the current word using _pext_u64 and merges them with the partial value (pt_val) from the previous word to reconstruct the complete integer (lines 26). Moreover, the presence of five leading 1s (11111) in mval indicates an incomplete integer extending into the next word. This partial integer is extracted and stored in pt_val for subsequent decoding, and the shift_bits variable is updated to 35 (the five 7-byte sequences) to align the decoding process for the next iteration. This decrement in n by 1 (line 28) ensures the algorithm’s seamless management of cross-boundary integers, maintaining decoding continuity and precision.
Case 63, depicted in lines 30-32, addresses a binary representation of 111111, signaling an integer spanning three data blocks, with the current block as its mid-segment. This specific case requires extracting six bytes (42 bits) via _pext_u64, appending this data to the partial value to maintain integer continuity across blocks.
In addition to the representative cases discussed above, our algorithm includes a comprehensive set of cases to handle various varint encoding patterns efficiently. These cases, omitted for brevity in the pseudocode (line 33), cover a wide range of scenarios, ensuring optimized decoding performance for different integer distributions.
The seamless handling of cross-boundary integers is a key strength of our approach. By maintaining state variables like shift_bits and pt_val, we ensure the continuity and precision of the decoding process across word boundaries. The decrement of n by the appropriate count in each case guarantees that the algorithm keeps track of the remaining integers to decode, maintaining the correctness of the output. Through this case analysis, we showcase the adaptability and efficiency of our BMI2-based varint decoding algorithm. By leveraging the power of the _pext_u64 instruction and carefully designing the decoding logic for each case, we can handle a wide range of varint encoding patterns with minimal branching and maximum performance.